OpenStudy (anonymous):

I really don't understand this problem (how to do it) at all, could someone please show and explain it to me? Given vectors u=-5i + 9j and v=4i+9j Find 2u-10v in terms of unit vectors i & j. (Sorry I can't make the vector signals)

OpenStudy (anonymous):

So...2u=-10i+18j and -10v=-40i-90j. 2u-10v=-10i+18j-40i-90j=-50i-62j. That's it i guess.

OpenStudy (anonymous):

You have to separately calculate the i and the j on the both: x axis and y axis.

OpenStudy (anonymous):

The answers are in multiple choice A. -71i - 51j B. -50i - 72j C. -77i - 51j D. -47i - 51j I don't know how to do that.

OpenStudy (anonymous):

ummm...yeah...my fault...90-18=72...not 62. So it's B.

OpenStudy (anonymous):

Alright, thank you very much.